# Evaluating LLMs’ Mathematical and Coding Competency through Ontology-guided Interventions

### v4

@@ I.1 Principles behind Ontology
Variable Response: Replace one value inside the question with a variable and answer with the variable included. Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has X boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. What is the total inner volume of all the boxes as a function of X? Code Execution: Given Docstring requirement, and specific input parameter, find the output for the function without writing any code. Changed from G.1: Find the output of the following function description, if the input is:string = “Hello World&7”
@@
Values Relationship: Identify the relationship between input values or parameters if the output or the final answer is given. Variable Relationship: Replace a pair of values inside the question with variables. After answering the original question, the variable forms a relationship. Query that relationship. Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has X boxes. Each box is Y inches by 6 inches by 4 inches. The walls are 1 inch thick. If the total inner volume of all the boxes is 72 cubic inches, then find the equation that relates X and Y? Parameter Relationship: Given the output of the function, categorize the possible groups of inputs parameters into the question. Changed from G.2: If the below program output integer 7. What is the relationship between a and b
@@
Variable Group: Change a group of several input values or parameters to variables. Variable Scaling: After answering the question, change the query to: if certain factual numbers in the question is scaled up by x, how will the final answer change? Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. Now, the number of boxes, the box outer dimensions, and the wall thickness are all scaled up by a factor of X. What is the total inner volume of all the boxes as a function of X? Variable Substitution: Change one or more variables inside the docstring to input parameters. Changed from G.1:
@@
Backward Reasoning: Reverse the reasoning process, reason from how to reach input from output. Variable Adaptation: If the answer to the question add or subtract by a certain amount x, pick one value inside the Information and ask how it should change if other values are kept the same. Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. If the total inner volume of all 3 boxes increases by a certain variable X, how should the thickness of the walls adjust correspondingly if the number of boxes and the external dimensions of each box stay the same? Write the answer as a function of X. Reverse Engineering: Change the Docstring, Function Header, and Examples to find the function that can reverse engineer the original function. Specifically, mapping the output back to its original inupt. Changed from G.1: Create a function that reverses the following function’s process, effectively transforming its output back into the original input
@@
What If: What would the outcome be if X had happened instead of Y, given the same initial conditions and context. WhatIf Question: First mask some number of values inside the question and answer the original question. What if we change one value inside the question, how will the final answer change? (The final answer should not have variables included as the masked value could be solved given the final answer.) Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches in width by 6 inches in length and a few inches in height. The walls are 1 inch thick. The total inner volume of all the boxes combined is 72 cubic inches. Now, if the thickness of the walls is half of its original thickness, then what will be the total inner volume? WhatIf Code: WhatIf the code structure or input value is changed, and some condition is masked. Changed from G.1: Find the output of the ‘changed_function‘, if the input is the same.
@@
Solve Value: Mask one variable’s value inside question, given answer, infer the masked value. Solve X: Replace one value inside the question with X and solve for X. Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has X boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. If the total inner volume of all 3 boxes is 72 cubic inches what is the value for X? Solve Input: Determine the input value of the function, based on the known output value. Changed from G.1 What is input to the following function, if the output is: "hELLO 9"
@@
Identify Range: Find what are possible constraint on the values. Variable Range: Replace one value with variable, and change the query to find the possible range of values based on the question. Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is X inches by 6 inches by 4 inches. The walls are 1 inch thick. Suppose we want to find out the total inner volume of all the boxes. What are the possible ranges of values of variable X based on the given information? Parameter Range: Identify what are the constraint on the input parameter, or what is the range of output parameter if input parameter is contraint to take certain value. Changed from G.3: If all the item inside the input list is smaller than 1, what are the constraints on the output from this function below?
@@ I.2 Concept Analysis
Identify Assumption: Identify one hidden commonsense assumption in the question that requires the answer to be answerable. Changed from F.3: {mdframed}[backgroundcolor=pink!20] You do not need to solve the question below, just identify one important hidden assumption that is required for the question to be answerable. Kylar went to the store to buy glasses for his new apartment. One glass costs $5, but every second glass costs only 60% of the price. Kylar wants to buy 16 glasses. How much does he need to pay for them? Test Case: List different boundary test cases that is valid for the input of the question. Changed from G.1: Provide input parameters for the test cases of the specified coding problem. These parameters should encompass boundary conditions within the scope defined by the function’s requirements specification, and avoid scenarios that fall outside of these requirements.
@@
Missing Info: Mask or delete an important piece of information and ask what additional information is needed to make the question answerable. Changed from F.2: {mdframed}[backgroundcolor=pink!20] John owns 3 boxes, each measuring 5 inches by 6 inches by 4 inches. Each box also had inner walls with certain non-zero thicknesses. Suppose you want to find out the total inner volume of all the boxes. What information is missing to calculate that? Incomplete Answer: Given the question, mask partial answer of the original, the model need to infer the missing lines based on the context. Changed from G.1: Complete the function below by predicting what is inside the masked code paragraph
Question Formulation: Formulate the question based on its answer. Question Formulation - (Math): Formulate a question to the chain of thought gold answer. Changed from F.1: {mdframed}[backgroundcolor=pink!20] Formulate a math application question that requires the following mathematical structure (calculations): 5000*(2.5/100) = $125 8000*(1.2/100) = $96 $125 > $96 $125 Math Question: Ask potential structures of math application. Question Formulation - (Code): Formulate a concise coding requirement by looking at the function code. Changed from G.2: Write a concise code description for the following code of its functionality no more than 1 sentence.
Add Misinformation: Add a piece of distracting information that can mislead the answer. Introduce Distraction: Add a Potentially Distracting information that will not affect the answer to the question. Changed from F.1: {mdframed}[backgroundcolor=pink!20] A merchant is considering a decision between the following purchase plans: jewelry with a value of $5,000, a trip to Europe costing $7,000, or electronic gadgets worth $8,000. His financial advisor predicts that the jewelry market will increase by 2.5%, the travel market will stay relatively stable with little to no change, and the electronic gadgets market will rise by 1.2% within the same month. He recently also came into an inheritance of $20,000 that he doesn’t need to use right away. If the merchant’s goal is to maximize profit at the end of this month by making a purchase choice, how much profit would this be? Introduce Bias: Change the python header to describe another function requirement, and change all the examples demonstrations bias towards a specific output Changed from G.1
Solution Evaluation: This dimension focuses on the problem-solving process to get to the final answer and how to optimize it.
@@
Optimize Solution: Assess whether the current state is optimal or if improvements are necessary
@@
Info Necessity: Check If there is redundant information given in the question, if yes, identify the redundant information. Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. Suppose we want to find out the total inner volume of all 3 boxes. To solve this math question, is there a way to determine the total inner volume of all 3 boxes without calculating the inner volume of one box? Reduce Complexity: Assess whether the complexity of the current code be further reduced. Changed from G.3: Optimize the code below to more efficiently achive the same requirement specified in the docstring
@@
Step Functionality: Whether there are alternative answers that follow the constraint. Step Necessity: Whether there are any alternative solutions reasoning steps without calculating an specific intermediate value. Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. Suppose we want to find out the total inner volume of all 3 boxes. To solve this math question, is there a way to determine the total inner volume of all 3 boxes without calculating the inner volume of one box? Step Necessity: Provide one line of code inside the Python function, and explain the functionality of that line of code in the context of the whole solution. Changed from G.2: Explain what is the the line below the comment functionality?
@@
Theoretical Basis: Identify the theory or principles in solving the question in general. Theoretical Basis (Math): Identify the underlying arithmetic or algebraic rules (toolbox) that govern the solution to the question. Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. Assume you want to find out the total inner volume of all 3 boxes. Can you identify one underlying mathematical theory which is required to do that? Theoretical Basis (Code): Request explanation on essential python concepts required to solve the question, for example, related to python objects and programming skills. Changed from G.1: Please describe to me in simple terms, assuming I have no knowledge of programming. Your task isn’t to solve the coding problem itself, but rather to identify the programming concepts in Python that would be necessary to address the problem presented below.
@@
Cost Analysis: Analyze the computational cost regarding the solution. Solution Efficiency: Compare two solution plans on solving the question and evaluate which one uses less computation. Changed from F.2: {mdframed}[backgroundcolor=pink!20] Evaluate which solution plan is more efficient in solving the question?
@@
Plan 1: Calculate the volume of the outer dimensions for one box, calculate the volume of the material used for the walls for one box, subtract the latter from the former to find the inner volume of one box, and then multiply this by 3 for all boxes.
@@
Plan 2: Calculate the inner dimensions of a single box by subtracting twice the thickness of the walls from each outer dimension, then find the volume of this inner space and multiply by 3 for all boxes. Code Complexity: Analyze the time complexity and space complexity of the provided code solution. Changed from G.1 Analyze the time and space complexity regarding to input parameter string of the following function.
@@
Spot Error: In this dimension, deliberate errors are introduced into the question or in a provided example answer. The purpose is to see if the LLM can identify and rectify these errors. This tests the LLM’s error detection capabilities, which is crucial for reliability in practical applications.
@@
Seek Clarification: The question requires to be clarified first before answering
@@
Introduce Ambiguity: Introduce Ambiguity to the question implicitly by changing the original information, so that the question cannot be solved without clarification. Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has three 5x6x4 inch boxes. A particular side of each box have 1 inch thick walls. What does the total inner capacity of these boxes amount to? Example Requirement: Remove the coding requirement in the docstring, instead only provide examples as a coding requirement. The provided examples will define and demonstrate the expected behavior in various scenarios. Changed from G.1: Begin by analyzing the function’s behavior specified in the docstring to understand its pattern, and then proceed to code the function accordingly.
@@
Conditional Analysis: Based on different possible situations of the question, the answer should separately presented. Discuss Separately: Introduce new information containing variables or conditions that require the answer to be discussed separately based on conditions or variables. Changed from F.1: {mdframed}[backgroundcolor=pink!20] A merchant wants to make a choice of purchase between 2 purchase plans: jewelry worth $5,000 or electronic gadgets worth $8,000. His financial advisor speculates that the jewelry market will go up x% while the electronic gadgets market will rise 1.2% within the same month. If the merchant is looking to maximize profit at the end of this month by making a choice, how much profit would this be? Incomplete requirement: Left some condition unspecified in the docstring. Changed from G.1:
@@
Conflicting Information: Introduce a new piece of information that is conflicting with existing information. This will make the question unanswerable, so the if the LLM can spot the error without mentioning. Introduce Contradiction: Add a piece of contradicting information to the question and check if LLM can spot the problem. Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. Each box is also 8 inches in width. What is the total inner volume of all 3 boxes? Wrong Example: Include an example that is conflicting with the requirement specified in the docstring. Changed from G.1:
@@
surface Error: Introduce an obvious error that can be spot without reasoning. Value Uncommon: Change the values so that it seems wired or unusual by commonsense knowledge standards. Changed from F.2: {mdframed}[backgroundcolor=pink!20] Can you spot anything unusual for the following question? John has 3 boxes. Each box measures 50000 miles by 60000 miles by 40000 miles. The walls of the boxes are 100 miles thick. What is the total inner volume of all 3 boxes? Runtime Error: Introduce a piece of error that will cause a runtime error or syntax error in python. Changed from G.1: Debug the error in the following code
@@
Hidden Error: Introduce a hidden error that need logical reasoning to spot. Value Error: Change the values so that the question does not make sense. Changed from F.4: {mdframed}[backgroundcolor=pink!20] Vicki is planning a pop concert at her high school. The show will be 2 minutes. She is allowing each group 2 hours to get on stage, 6 hours to perform, and then 2 hours to exit the stage. If she allows a 10-hour intermission, how many groups can perform in the concert? Value Error: Introduce the change in the code that will cause a Value Error in python. Changed from G.1
I.3 Representational Perturbation – Format Change

### v5

@@ I.1 Principles behind Ontology
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has X boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. What is the total inner volume of all the boxes as a function of X? Code Execution: Given Docstring requirement, and specific input parameter, find the output for the function without writing any code. Changed from G.1: Find the output of the following function description, if the input is:string = “Hello World&7”1def flip_case(string: str) -> str:2 """For a given string, flip lowercase characters to uppercase and uppercase to lowercase.""" -
G20.
Value Relationship: Identify the relationship between input values or parameters if the output or the final answer is given.
Variable Relationship: Replace a pair of values inside the question with variables. After answering the original question, the variable forms a relationship. Query that relationship.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has X boxes. Each box is Y inches by 6 inches by 4 inches. The walls are 1 inch thick. If the total inner volume of all the boxes is 72 cubic inches, then find the equation that relates X and Y? Parameter Relationship: Given the output of the function, categorize the possible groups of inputs parameters into the question. Changed from G.2: If the below program output integer 7. What is the relationship between a and b -
G21.
Variable Group: Change a group of several input values or parameters to variables.
Variable Scaling: After answering the question, change the query to: if certain factual numbers in the question is scaled up by x, how will the final answer change?
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. Now, the number of boxes, the box outer dimensions, and the wall thickness are all scaled up by a factor of X. What is the total inner volume of all the boxes as a function of X? Variable Substitution: Change one or more variables inside the docstring to input parameters. Changed from G.1:1def flip_case(string: str, specific_value: str) -> str:23 """"""For a given string and specific value, flip the specific value from lowercase to uppercase or uppercase to lowercase. The function will only flip the case of the specific value in the string.4 >>> flip_case(’Hello’, ’h’)5 ’hello’6 """" -
G22.
Backward Reasoning: Reverse the reasoning process, reason from how to reach input from output.
Variable Adaptation: If the answer to the question add or subtract by a certain amount x, pick one value inside the Information and ask how it should change if other values are kept the same.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. If the total inner volume of all 3 boxes increases by a certain variable X, how should the thickness of the walls adjust correspondingly if the number of boxes and the external dimensions of each box stay the same? Write the answer as a function of X. Reverse Engineering: Change the Docstring, Function Header, and Examples to find the function that can reverse engineer the original function. Specifically, mapping the output back to its original inupt. Changed from G.1: Create a function that reverses the following function’s process, effectively transforming its output back into the original input -
G23.
Counterfactual: What would the outcome be if X had happened instead of Y, given the same initial conditions and context.
WhatIf Question: First mask some number of values inside the question and answer the original question. What if we change one value inside the question, how will the final answer change? (The final answer should not have variables included as the masked value could be solved given the final answer.)
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches in width by 6 inches in length and a few inches in height. The walls are 1 inch thick. The total inner volume of all the boxes combined is 72 cubic inches. Now, if the thickness of the walls is half of its original thickness, then what will be the total inner volume? WhatIf Code: WhatIf the code structure or input value is changed, and some condition is masked. Changed from G.1: Find the output of the ‘changed_function‘, if the input is the same.1We know that if we input masked_input to the ‘original_function‘, the output is following:2>>> original_function(masked_input)3’hELLO’45Here is the ‘original_function‘:6def original_function(string: str) -> str:7 return string.swapcase()89Here is the ‘changed_function‘:10def changed_function(string: str) -> str:11 return string.swapcase()[::-1]1213What will be the output for ‘changed_function(masked_input)‘"" -
G24.
Solve Value: Mask one variable’s value inside question, given answer, infer the masked value.
Solve X: Replace one value inside the question with X and solve for X.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has X boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. If the total inner volume of all 3 boxes is 72 cubic inches what is the value for X? Solve Input: Determine the input value of the function, based on the known output value.
Changed from G.1 What is input to the following function, if the output is: "hELLO 9" -
G25.
Identify Range: Find what are possible constraint on the values.
Variable Range: Replace one value with variable, and change the query to find the possible range of values based on the question.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is X inches by 6 inches by 4 inches. The walls are 1 inch thick. Suppose we want to find out the total inner volume of all the boxes. What are the possible ranges of values of variable X based on the given information? Parameter Range: Identify what are the constraint on the input parameter, or what is the range of output parameter if input parameter is contraint to take certain value.
Changed from G.3: If all the item inside the input list is smaller than 1, what are the constraints on the output from this function below?
@@ I.2 Concept Analysis
G23.
Inherent Premise: Identify the underlying premise of the question.
Identify Assumption: Identify one hidden commonsense assumption in the question that requires the answer to be answerable.
Changed from F.3: {mdframed}[backgroundcolor=pink!20] You do not need to solve the question below, just identify one important hidden assumption that is required for the question to be answerable. Kylar went to the store to buy glasses for his new apartment. One glass costs $5, but every second glass costs only 60% of the price. Kylar wants to buy 16 glasses. How much does he need to pay for them? Test Case: List different boundary test cases that is valid for the input of the question. Changed from G.1: Provide input parameters for the test cases of the specified coding problem. These parameters should encompass boundary conditions within the scope defined by the function’s requirements specification, and avoid scenarios that fall outside of these requirements.1def flip_case(string: str) -> str:2 """For a given string, flip lowercase characters to uppercase and uppercase to lowercase.3 """ -
G24.
@@
Info Sufficiency: Mask or delete an important piece of information and ask what additional information is needed to make the question answerable.
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John owns 3 boxes, each measuring 5 inches by 6 inches by 4 inches. Each box also had inner walls with certain non-zero thicknesses. Suppose you want to find out the total inner volume of all the boxes. What information is missing to calculate that? Incomplete Answer: Given the question, mask partial answer of the original, the model need to infer the missing lines based on the context. Changed from G.1: Complete the function below by predicting what is inside the masked code paragraph1def flip_case(string: str) -> str:2 """For a given string, flip lowercase characters to uppercase and uppercase to lowercase.3 >>> flip_case(’Hello’)4 ’hELLO’5 """6 [masked code paragraph]7 if char.isupper():8 result += char.lower()9 else:10 result += char.upper()11 return result -
G25.
Question Formulation: Formulate the question based on its answer.
Question Formulation - (Math): Formulate a question to the chain of thought gold answer.
Changed from F.1: {mdframed}[backgroundcolor=pink!20] Formulate a math application question that requires the following mathematical structure (calculations): 5000*(2.5/100) = $125 8000*(1.2/100) = $96 $125 > $96 $125 Math Question: Ask potential structures of math application. Question Formulation - (Code): Formulate a concise coding requirement by looking at the function code.
Changed from G.2: Write a concise code description for the following code of its functionality no more than 1 sentence. -
G26.
Add Misinformation: Add a piece of distracting information that can mislead the answer.
Introduce Distraction: Add a Potentially Distracting information that will not affect the answer to the question.
Changed from F.1: {mdframed}[backgroundcolor=pink!20] A merchant is considering a decision between the following purchase plans: jewelry with a value of $5,000, a trip to Europe costing $7,000, or electronic gadgets worth $8,000. His financial advisor predicts that the jewelry market will increase by 2.5%, the travel market will stay relatively stable with little to no change, and the electronic gadgets market will rise by 1.2% within the same month. He recently also came into an inheritance of $20,000 that he doesn’t need to use right away. If the merchant’s goal is to maximize profit at the end of this month by making a purchase choice, how much profit would this be? Introduce Bias: Change the python header to describe another function requirement, and change all the examples demonstrations bias towards a specific output Changed from G.11def uppercase(string: str) -> str:2 """For a given string, flip lowercase characters to uppercase and uppercase to lowercase.3 >>> flip_case(’hello’)4 ’HELLO’5 """
G23.
Solution Understanding: This dimension focuses on the problem-solving process to get to the final answer and how to optimize it.
@@
Optimize Solution: Assess whether the current state is optimal or if improvements are necessary.
Info Necessity: Check If there is redundant information given in the question, if yes, identify the redundant information.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. Suppose we want to find out the total inner volume of all 3 boxes. To solve this math question, is there a way to determine the total inner volume of all 3 boxes without calculating the inner volume of one box? Reduce Complexity: Assess whether the complexity of the current code be further reduced.
Changed from G.3: Optimize the code below to more efficiently achive the same requirement specified in the docstring1def derivative_polynomial(coefficients, derivative=None, index=0):2 """3 This function calculates the derivative of a polynomial using recursion.4 coefficients: List of coefficients of the polynomial.5 derivative: List to store the coefficients of the derivative. Initially None.6 index: Current index in the coefficients list.78 The base case of the recursion is when the index is equal to the length of the coefficients list.9 """10 if index > 0:11 derivative_coefficient = index * coefficients[index]12 derivative.append(derivative_coefficient)13 return derivative_polynomial(coefficients, derivative, index + 1) -
G28.
Step Functionality: Whether there are alternative answers that follow the constraint.
Step Necessity: Whether there are any alternative solutions reasoning steps without calculating an specific intermediate value.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. Suppose we want to find out the total inner volume of all 3 boxes. To solve this math question, is there a way to determine the total inner volume of all 3 boxes without calculating the inner volume of one box? Step Necessity: Provide one line of code inside the Python function, and explain the functionality of that line of code in the context of the whole solution.
Changed from G.2: Explain what is the the line below the comment functionality?1def greatest_common_divisor(a: int, b: int) -> int:23 """ Return a greatest common divisor of two integers a and b4 >>> greatest_common_divisor(3, 5)5 16 >>> greatest_common_divisor(25, 15)7 58 """9 while b:10 a, b = b, a % b11 # What is the functionality of ‘abs()‘12 return abs(a) -
G29.
Theoretical Basis: Identify the theory or principles in solving the question in general.
Theoretical Basis (Math): Identify the underlying arithmetic or algebraic rules (toolbox) that govern the solution to the question.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. Assume you want to find out the total inner volume of all 3 boxes. Can you identify one underlying mathematical theory which is required to do that? Theoretical Basis (Code): Request explanation on essential python concepts required to solve the question, for example, related to python objects and programming skills.
Changed from G.1: Please describe to me in simple terms, assuming I have no knowledge of programming. Your task isn’t to solve the coding problem itself, but rather to identify the programming concepts in Python that would be necessary to address the problem presented below.1def flip_case(string: str) -> str:2 """For a given string, flip lowercase characters to uppercase and uppercase to lowercase.3 >>> flip_case(’Hello’)4 ’hELLO’5 """ -
G30.
Cost Analysis: Analyze the computational cost regarding the solution.
@@
Solution Efficiency: Compare two solution plans on solving the question and evaluate which one uses less computation.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] Evaluate which solution plan is more efficient in solving the question?Plan 1: Calculate the volume of the outer dimensions for one box, calculate the volume of the material used for the walls for one box, subtract the latter from the former to find the inner volume of one box, and then multiply this by 3 for all boxes.
@@
Plan 2: Calculate the inner dimensions of a single box by subtracting twice the thickness of the walls from each outer dimension, then find the volume of this inner space and multiply by 3 for all boxes. Code Complexity: Analyze the time complexity and space complexity of the provided code solution.
Changed from G.1 Analyze the time and space complexity regarding to input parameter string of the following function.1def flip_case(string: str) -> str:2 """For a given string, flip lowercase characters to uppercase and uppercase to lowercase.3 >>> flip_case(’Hello’)4 ’hELLO’5 """
G27.
@@
Critical Thinking: In this dimension, deliberate errors are introduced into the question or in a provided example answer. The purpose is to see if the LLM can identify and rectify these errors. This tests the LLM’s error detection capabilities, which is crucial for reliability in practical applications.
@@
Seek Clarification: The question requires to be clarified first before answering.
Introduce Ambiguity: Introduce Ambiguity to the question implicitly by changing the original information, so that the question cannot be solved without clarification.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has three 5x6x4 inch boxes. A particular side of each box have 1 inch thick walls. What does the total inner capacity of these boxes amount to? Example Requirement: Remove the coding requirement in the docstring, instead only provide examples as a coding requirement. The provided examples will define and demonstrate the expected behavior in various scenarios. Changed from G.1: Begin by analyzing the function’s behavior specified in the docstring to understand its pattern, and then proceed to code the function accordingly.1def flip_case(string: str) -> str:2 """3 function(’Hello’) == ’hELLO’4 function(’Python 3.8’) == ’pYTHON 3.8’5 function(’123abcXYZ’) == ’123ABCxyz’6 function(’MixedCASE123’) == ’mIXEDcase123’7 function(’ALLUPPERCASE’) == ’alluppercase’8 """ -
G32.
Conditional Analysis: Based on different possible situations of the question, the answer should separately presented.
Discuss Separately: Introduce new information containing variables or conditions that require the answer to be discussed separately based on conditions or variables.
@@
Changed from F.1: {mdframed}[backgroundcolor=pink!20] A merchant wants to make a choice of purchase between 2 purchase plans: jewelry worth $5,000 or electronic gadgets worth $8,000. His financial advisor speculates that the jewelry market will go up x% while the electronic gadgets market will rise 1.2% within the same month. If the merchant is looking to maximize profit at the end of this month by making a choice, how much profit would this be? Incomplete requirement: Left some condition unspecified in the docstring. Changed from G.1:1def flip_case(ch: str) -> str:23 """For a given string, all the letters inside the string should be changed. flip lowercase characters to uppercase.4 >>> flip_case(’h’)5 ’H’6 """ -
G33.
Conflicting Information: Introduce a new piece of information that is conflicting with existing information. This will make the question unanswerable, so the if the LLM can spot the error without mentioning.
Introduce Contradiction: Add a piece of contradicting information to the question and check if LLM can spot the problem.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. Each box is also 8 inches in width. What is the total inner volume of all 3 boxes? Wrong Example: Include an example that is conflicting with the requirement specified in the docstring. Changed from G.1:1def flip_case(string: str) -> str:2 """"""For a given string, flip lowercase characters to uppercase and uppercase to lowercase.3 >>> flip_case(’Hello’)4 ’hello’5 """" -
G34.
Surface Error: Introduce an obvious error that can be spotted without reasoning.
Value Uncommon: Change the values so that it seems weird or unusual by commonsense knowledge standards.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] Can you spot anything unusual in the following question? John has 3 boxes. Each box measures 50000 miles by 60000 miles by 40000 miles. The walls of the boxes are 100 miles thick. What is the total inner volume of all 3 boxes? Runtime Error: Introduce a piece of error that will cause a runtime error or syntax error in python. Changed from G.1: Debug the error in the following code1def flip_case(string, str) -> str:2 """For a given string, flip lowercase characters to uppercase and uppercase to lowercase.3 >>> flip_case(’Hello’)4 ’hELLO’5 """6 return string.swapcase() -
G35.
Hidden Error: Introduce a hidden error that need logical reasoning to spot.
Value Error: Change the values so that the question does not make sense.
@@
Changed from F.4: {mdframed}[backgroundcolor=pink!20] Vicki is planning a pop concert at her high school. The show will be 2 minutes. She is allowing each group 2 hours to get on stage, 6 hours to perform, and then 2 hours to exit the stage. If she allows a 10-hour intermission, how many groups can perform in the concert? Logical Error: Introduce the change in the code that will cause a Logical Error in Python. Changed from G.11def flip_case(string: str) -> str:2 """For a given string, flip lowercase characters to uppercase and uppercase to lowercase.3 >>> flip_case(’Hello’)4 ’hELLO’5 """6 string = list(string.swapcase())7 return string
G31.
I.3 Question Format